The sum of two counting numbers is 45. One of the number exceeds the other by 7. What are the two numbers?
Answer by ikleyn(52781)

.

    x + (x+7) = 45


      2x      = 45 - 7 = 38


       x               = 38/2


ANSWER.  The numbers are 19 and 26.

Solved.